# Pulsar Timing Noise from Brownian Motion of the Sun

### v2

@@ 2 Brownian Motion of the Sun
Previous analysis of PTA data was based on the Solar system ephemeris calculated from well-studied objects in the Solar system (see Agazie et al. (2023c) and Vallisneri et al. (2020), and references therein). The Solar system ephemeris model includes 343 asteroids (Park et al., 2021), but according to the cumulative number versus diameter in Davis et al. (2002), there are many more than 343 Main belt asteroids below a mass scale of $\sim 10^{21}~{}{\rm g}$.
@@
Here, I focus on the large population of unmodeled asteroids ($\lesssim 80$km in diameter) as a source of residual noise in correlated pulsar timing. Squaring both sides of equation (1), taking a statistical average (denoted hereafter by angular brackets, following the ergodic theorem) and ignoring velocity correlations among the many asteroids, I get the dispersion in random velocity fluctuations of the Sun as a result of its stochastic gravitational interactions with the cluster of unmodeled asteroids,
@@
$\delta v\equiv\langle|v_{\odot}^{2}|\rangle^{1/2}=7\times 10^{-5}~{}{\rm cm~{}s^{-1}}\left({N_{a}\over 10^{3}}\right)^{1/2}\left({m_{a}\over 10^{21}~{}{\rm g}}\right)\left({\langle|v_{a}^{2}|\rangle^{1/2}\over 20~{}{\rm km~{}s^{-1}}}\right)~{}.$ (3)
@@
$h_{c}\sim\left({\delta v\over c}\right)\sim 2\times 10^{-15}\left({N_{a}\over 10^{3}}\right)^{1/2}\left({m_{a}\over 10^{21}~{}{\rm g}}\right)\left({\langle|v_{a}^{2}|\rangle^{1/2}\over 20~{}{\rm km~{}s^{-1}}}\right)~{},$ (4)
@@ 3 Discussion
An increase in the number of Main-belt asteroids included in the Solar system ephemeris (Park et al., 2021) and future observations of these asteroids with the Large Survey of Space and Time (LSST) of the Vera C. Rubin Observatory in Chile (Schwamb et al., 2023), would allow better modeling of the noise they introduce to PTAs.
@@ Acknowledgements
I thank Matt Holman and Peter Veres for helpful comments. This work was supported in part by Harvard’s Black Hole Initiative, which is funded by grants from JFT and GBMF.

### v3

@@ 2 Brownian Motion of the Sun
Previous analysis of PTA data was based on the Solar system ephemeris calculated from well-studied objects in the Solar system (see Agazie et al. (2023c) and Vallisneri et al. (2020), and references therein). The Solar system ephemeris model includes 343 asteroids (Park et al., 2021), but according to the cumulative asteroid number versus diameter in Davis et al. (2002), there are many more than 343 Main belt asteroids below a mass scale of $\sim 10^{21}~{}{\rm g}$.
@@
Here, I focus on the large population of unmodeled asteroids ($\lesssim 80$km in diameter) as a source of residual noise in correlated pulsar timing (see also, Fedderke et al. (2021)). Squaring both sides of equation (1), taking a statistical average (denoted hereafter by angular brackets, following the ergodic theorem) and ignoring velocity correlations among the many asteroids, I get the dispersion in random velocity fluctuations of the Sun as a result of its stochastic gravitational interactions with the cluster of unmodeled asteroids,
@@
$\delta v\equiv\langle|v_{\odot}^{2}|\rangle^{1/2}=3.2\times 10^{-5}~{}{\rm cm~{}s^{-1}}\left({N_{a}\over 10^{3}}\right)^{1/2}\left({m_{a}\over 10^{21}~{}{\rm g}}\right)\left({\langle|v_{a}^{2}|\rangle^{1/2}\over 20~{}{\rm km~{}s^{-1}}}\right)~{}.$ (3)
@@
$h_{c}\sim\left({\delta v\over c}\right)\sim 10^{-15}\left({N_{a}\over 10^{3}}\right)^{1/2}\left({m_{a}\over 10^{21}~{}{\rm g}}\right)\left({\langle|v_{a}^{2}|\rangle^{1/2}\over 20~{}{\rm km~{}s^{-1}}}\right)~{},$ (4)
@@ 3 Discussion
When squaring the right-hand-side of equation (1), we ignored correlations between the velocities of different unmodeled asteroids. Any such correlations as a result of the phase-space clustering of asteroids due to planets, like Jupiter’s clusters of Trojans, Greeks and Hildas, would enhance the expected noise. In that case, the mass of a perturber will be that of a cluster of asteroids rather than a single asteroid. For example, there are more than 100 Jupiter’s Trojans with diameters in the range 40-80km, exceeding $10^{22}$g in total mass. The additional effect of clustering is an enhancement of the noise amplitude in equation (3) by using the cluster mass and number in place of $m_{a}$ and $N_{a}$, which cannot exceed an extra factor of $\sqrt{N_{a}}$ altogether - the upper limit associated with placing all asteroids in a single cluster.
@@
An increase in the number of Main-belt asteroids included in the Solar system ephemeris (Park et al., 2021) and future observations of missed asteroids with the Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory in Chile (Schwamb et al., 2023), would allow better modeling of the noise they introduce to PTAs.
@@ Acknowledgements
I thank Michael Fedderke for correcting a typo and sharing insights, and Matt Holman and Peter Veres for helpful comments. This work was supported in part by Harvard’s Black Hole Initiative, which is funded by grants from JFT and GBMF.